OpenStudy (yinshy):

Emily invested $6500 at 3% interest compounded annually. How much interest will she earn in 5 years? A. $975.00 B. $1035.28 C. $7475.00 D. $7535.28

OpenStudy (greatlife44):

Again we need to use the compound interest formula: \[A = P(1+\frac{ r}{ n })^{nt}\]

OpenStudy (yinshy):

C?

OpenStudy (greatlife44):

\[6,500(1+\frac{ 0.03 }{ 1 })^{1*5} = 7535.28\]